# CAN QUANTUM THEORY BE UNDERPINNED BY A NON-LOCAL HIDDEN VARIABLE THEORY ?

### v1

@@ B J Dalton
1 Abstract
In this paper we consider the description by a Bell non-local hidden variable theory of bipartite quantum states with two observables per sub-system. We derive Bell inequalities of the Collins-Gisin.-Liden-Massar-Popescu type which involve combinations of the probabilities of related outcomes for measurements for the four pairs of sub-system observables. It is shown that the corresponding quantum theory expressions violate the Bell inequalities in the case of the maximally entangled state of the bipartitite system. This shows that quantum theory can not be underpinned by a non-local hidden variable theory. So as local hidden variable theory has already been shown to conflict with quantum theory, it follows that quantum theory can not be understood in terms of any hidden variable theory (which Einstein originally had expected).
@@ 2 INTRODUCTION
This paper deals with the issue of whether quantum theory needs to be completed or underpinned by a non-local hidden variable theory (NLHVT), an issue that originates from the 1930’s papers of Einstein [1] and Schrodinger ([2], [3]) and the 1960’s paper of Bell ([4]). The key word is non-local,as it is now widely accepted that quantum theory is not underpinned by a local hidden variable theory.
@@ 2.1 Quantum Paradoxes
The 1930’s papers identified two paradoxical features arizing from the Copenhagen interpretation of quantum theory (QTHY), even for the simplest case of bipartite systems. The first was macroscopic entanglement – such as in the Schrodinger cat experiment, wherein states could exist in bipartite systems in which a sub-system could simultaneously occupy two macroscopically distinct states. Here the system state was an entangled state of a cat and a radioactive atom, involving both the cat being dead and the cat being alive. This conflicted with common sense - the cat should be either dead or alive. The second was EPR steering - measuring the value for an observable chosen for one sub-system $A$ sometimes could instantaneously affect the outcome for measuring the value of an observable in a second sub-system $B$ that was spatially well-separated. This contradicted relativity and for a system involving two particles in a state with a well-defined position difference, also led to a violation of the Heisenberg U Principle in terms of evidently being able to assign precise values to the position and momentum of one of the two particles. So maybe the particle really does have both a definite position and momentum - and the cat is either dead or alive, whilst the probababilistic feature of quantum theory is just similar to that in classical statistical mechanics ?
@@ 2.2 Quantum Theory Incompleteness ?
Thus arose Einstein’s question: [1]. Does quantum theory (though agreeing with experiment) require completion via an underlying theory that determines the real values of physical quantities - such as the position and momentum (for the particle) or being alive or being dead (for the cat), even if these are hidden within the probability distribution ?  Such underlying theories are generally referred to as hidden variable theories.(HVT ).
2.3 Local Hidden Variable Theory and Bell Non-Locality
@@
LHVT provides predictions for inequalities involving results based on mean values or the probabilities of joint measurement outcomes of pairs of observables from the two sub-systems - generically referred to as Bell inequalities. We can then test whether the inequality predictions based on LHVT are consistent both with quantum theory predictions and with experiment. As described below, some Bell inequality violations show LHVT fails. So does this mean that quantum theory could be underpinned or completed via a non-local HVT (NLHVT) ?
@@
In this paper a non-local hidden variable theory will be presented, and Bell inequalities obtained based on this NLHVT. These inequalities are of the Collins, Gisin, Linden, Massar Popescu (CGLMP) type (see [5], [6]). It will then be shown that in a maximally entangled state in a bipartite system clear violations of these Bell inequalities occur, thus showing that it is not posible to underpin or complete quantum theory via a non-local hidden variable theory. As well as being an important quantum foundations issue, there are also implications also for quantum technology, since systems for secure communications, quantum information processing, etc depend on there being different outcomes predicted for quantum theory and for hidden variable theory.
It should be noted that previous to the work of Bell [4], a different kind of non-local hidden variable theory was introduced by Bohm [7] in the 1950’s, based on earlier work by de Broglie . The approach involves deterministic particle trajectories dependent on a so-called quantum  potential determined (in the non-relativistic version) from the multi-particle Schrodinger wave function for the whole universe. The particle positions act like hidden variables and non-locality refers to the feature that the velocity of any one particle depends on the positions of all the other particles, no matter where their relative locations may be. Interest in Bohmian mechanics and its extensions still continues [8].
@@ 2.4 Plan of Paper
First however, we need to outline the key equations describing both LHVT and NLHVT and present a well-known example of a Bell inequality that is inconsistent with quantum theory. Next we point out how Bell inequalities are also relevant to discussing quantum entanglement and EPR steering based on  a classification of Bell local states into three categories involving so-called local hidden quantum states [9]. The material outlining this is in Appendix A. We then introduce the so-called CGLMP Bell inequalities [5] and set out the relevant equations for the inequalities involving the quantity $S$, the observables and the test quantum state, both for the LHVT and NLHVT cases. The next section covers the basic matrix equation method for determining the requirements for possible Bell inequalities of the form $S\leq 4,S\leq 3,S\leq 2,S\leq 1$ in the NLHVT situation  The Bell test requiremennts in each of these cases are set out, with the detailed derivations being presented in Appendix B. The last part of this section then sets out the quantum theory expressions for $S$. A comparison of the quantum theory numerical results for $S$ with the NLHVT results then follows. The paper ends with a Summary and Conclusions.
@@ 3 QTHY & HVT BASIC EQUATIONS
QTHY and LHVT joint probabilities $P(\alpha,\beta|\Omega_{A},\Omega_{B})$ for measurements of observables $\Omega_{A}$, $\Omega_{B}$ for sub-systems $A$, $B$ with outcomes $\alpha$, $\beta$ are:
@@
For QTHY $\widehat{\Pi}_{\alpha}^{A}$ is the projector onto the space of eigenvectors for operator $\widehat{\Omega}_{A}$ with eigenvalue $\alpha$. $\widehat{\rho}$ is density opr specifying the preparation process $c$ of quantum state. For LHVT, hidden variables $\lambda$ determined with probability $P(\lambda|c)$ for preparation process $c$.
In the probabilistic form of LHVT the hidden variables determine each local sub-system measurement probabilities  $P(\alpha|\Omega_{A},\lambda)$, $P(\beta|\Omega{}_{B},\lambda)$ for classical observables $\Omega_{A}$, $\Omega_{B}$ which  are combined using classical probability theory to determine joint measurement probabilities. Since the probabilities for all outcomes must add up to unity we have $\sum_{\alpha}P(\alpha|\Omega_{A},\lambda)=1$  and $\sum_{\beta}P(\beta|\Omega_{B},\lambda)=1$.
In the probabilistic form of NLHVT the hidden variables determine the combined sub-system measurement probabilities  $P(\alpha,\beta|\Omega_{A},\Omega_{B},\lambda)$, for classical observables $\Omega_{A}$, $\Omega_{B}$  In the non–local hidden variable theory there is no separate (or local) HVT probability for each sub-system.
@@
Again, the sum of the probabilities for all outcomes must be unity,so $\sum_{\alpha,\beta}P(\alpha,\beta|\Omega_{A},\Omega_{B},\lambda)=1$.
@@ 3.1 Joint Measurement Mean Values
where $\left\langle\Omega_{A}(\lambda)\right\rangle=\sum_{\alpha}\alpha P(\alpha|\Omega_{A},\lambda)$ is LHVT mean value for measurement of $\Omega_{A}$ when hidden variables are $\lambda$.
@@
3.2 Bell Inequality Test for Bell-Locality
@@
$S$ $=$ $|\left\langle\Omega_{A1}\otimes\Omega_{B1}\right\rangle_{LHVT}+\left\langle\Omega_{A1}\otimes\Omega_{B2}\right\rangle_{LHVT}+\left\langle\Omega_{A2}\otimes\Omega_{B1}\right\rangle_{LHVT}-\left\langle\Omega_{A2}\otimes\Omega_{B2}\right\rangle_{LHVT}\,|\,\leq 2$
This is violated for singlet state of two spin $1/2$ systems (spin operators in units /2Planck-constant-over-$\hbar/2$).
@@
for $\Omega_{A1}=S_{z}^{A}$, $\Omega_{A2}=S_{x}^{A}$, $\Omega_{B1}=-(S_{x}^{B}+S_{z}^{B})/\sqrt{2}$, $\Omega_{B2}=+(S_{x}^{B}-S_{z}^{B})/\sqrt{2}$, find $S=2\sqrt{2}$.
3.3 Local Hidden Variable Theory Refuted

### v2

@@ B J Dalton
0.1 Abstract
In this paper we consider the description by a general Bell-type non-local hidden variable theory of bipartite quantum states with two observables per sub-system. We derive Bell inequalities of the Collins-Gisin.-Liden-Massar-Popescu type which involve combinations of the probabilities of related outcomes for measurements for the four pairs of sub-system observables. It is shown that the corresponding quantum theory expressions violate the Bell inequalities in the case of the maximally entangled state of the bipartitite system. The CHSH Bell inequality is also derived from this general CGLMP Bell-type non-local hidden variable theory. This shows that quantum theory can not be underpinned by a Bell-type non-local hidden variable theory. So as a general Bell-type local hidden variable theory has already been shown to conflict with quantum theory, it follows that quantum theory can not be understood in terms of any CGLMP Bell-type hidden variable theory - local or non-local.
@@ 1 INTRODUCTION
This paper deals with the issue of whether quantum theory needs to be completed or underpinned by a non-local hidden variable theory (NLHVT), an issue that originates from the 1930’s papers of Einstein [1] and Schrodinger ([2], [3]) and the 1960’s paper of Bell ([5]). The key word is non-local,as it is now widely accepted that quantum theory is not underpinned by a local hidden variable theory (LHVT).
@@ 1.1 Quantum Paradoxes
The 1930’s papers identified two paradoxical features arizing from the Copenhagen interpretation of quantum theory (QTHY), even for the simplest case of bipartite systems. The first was macroscopic entanglement – such as in the Schrodinger cat experiment, wherein states could exist in bipartite systems in which a sub-system could simultaneously occupy two macroscopically distinct states. Here the system state was an entangled state of a cat and a radioactive atom, a quantum superposition state involving both the cat being dead and the cat being alive. This conflicted with common sense - the cat should be either dead or alive. The second was EPR steering - measuring the value for an observable chosen for one sub-system $A$ sometimes could instantaneously affect the outcome for measuring the value of an observable in a second sub-system $B$ that was spatially well-separated. This contradicted relativity and for a system involving two particles in a possible state with a well-defined position difference (and a well-defined momentum sum), also led to a violation of the Heisenberg Uncertainty Principle in terms of evidently being able to assign precise values to both the position and momentum of one of the two particles. So maybe the particle really does have both a definite position and momentum - and the cat is either dead or alive, whilst the probabilistic feature of quantum theory is just similar to that in classical statistical mechanics ?
@@ 1.2 Quantum Theory Incompleteness ?
Thus arose Einstein’s question: [1]. Does quantum theory (though agreeing with experiment) require completion via an underlying theory that determines the actual values of physical quantities (realism) - such as the position and momentum (for the particle) or being alive or being dead (for the cat), even if these are hidden within the probability distribution ?  Since the work of Bohm [4] and Bell [5] such underlying (or underpinning) theories are now generally referred to as hidden variable theories.(HVT ).
1.3 Local and Non-Local Hidden Variable Theory
@@
LHVT provides predictions for inequalities involving results based on mean values or the probabilities of joint measurement outcomes of pairs of observables from the two sub-systems - generically referred to as Bell inequalities. We can then test whether the inequality predictions based on LHVT are consistent both with quantum theory predictions and with experiment. As described below, some Bell inequality violations show LHVT fails. So does this mean that quantum theory should be underpinned or completed via a Bell-type non-local HVT (NLHVT) ?
@@
In this paper a general non-local hidden variable theory will be presented, and Bell inequalities obtained based on this NLHVT.We will show that the famous CHSH [6] Bell inequality can be established from NLHVT as well as from LHVT. We then consider Bell inequalities of the Collins, Gisin, Linden, Massar Popescu (CGLMP) type (see [7], [8]). It will then be shown that in a maximally entangled state in a bipartite system there are clear violations of these Bell inequalities, thus showing that it is not posible to underpin or complete quantum theory via this general Bell-type non-local hidden variable theory. As well as being an important quantum foundations issue, there are also implications also for quantum technology, since systems for secure communications, quantum information processing, etc depend on there being different outcomes predicted for quantum theory and for hidden variable theory.
1.4 Bohmian Mechanics
It should be noted that previous to the work of Bell [5], a different kind of non-local hidden variable theory was introduced by Bohm [4] in the 1950’s, based on earlier work by de Broglie (see [9] for a recent account). The approach involves deterministic particle trajectories dependent on a so-called quantum potential determined (in the non-relativistic version) from the multi-particle Schrodinger wave function for the system. The particle positions act like hidden variables and non-locality refers to the feature that the velocity of any one particle depends on the positions of all the other particles, no matter where their relative locations may be. This would violate special relativity, but as the related version of quantum theory is also non-relativistic this outcome is merely being consistent. As the derivation of the equations for Bohmian mechanics (BM) invokes the existence of the wave-function, assumes the correctness of the time-dependent Schrodinger equation (TDSE) and that the modulus squared of the wave function is the position probability density for the particles, it may be concluded that Bohmian mechanics is based on just the same equations as quantum theory but with some extra assumptions and ontological features (the deterministic trajectories) added. The particle trajectory feature is based on deriving a continuity equation for the position probabilty density and identifying the gradient of the phase $S$ of the wave-function (divided by the partticle mass) with the velocity for the deterministic particle trajectory - whose existence is now incorporated into the theory. This identification is based on a continuity equation for the position probability derived from equations for the phase and amplitude of the wave-function obtained from the TDSE. So the addition of deterministic trajectories complies with Einstein’s suggestion of completing quantum theory by adding new features - the deterministic trajectories. Whether the particle positions should be regarded as ”hidden variables” is a moot point. Unlike the hidden variables $\lambda$ introduced by Bell [5] (see below), the particle positions could in principle be observed, and they obey well defined Hamilton-Jacobi equations involving the quantum potential - in contrast to the unspecified equations for the $P(\lambda|c)$ and $P(\alpha|\Omega_{A},\lambda)$ in Bell’s approach. BM is rather like classical  physics and uses a more familiar mathematics based on calculus rather than the mathematics of linear vector spaces and Hermitian  operators. Many of the characteristic results in standard quantum theory can also be shown in BM. For example the component $L_{z}$ of the orbital  angular momentum of a particle is $xp_{y}-yp_{x}$, which becomes $x(\partial S/\partial y)-y(\partial S/\partial x)$ $=(\partial S/\partial\phi)$ from the particle trajectory assumption in which $p_{y}=mv_{y}$ etc.. As the stationary solution of the Schrodinger equation for a spherical potential involves a phase $S=\beta-Et+\hslash m\phi$ ($m$ an integer) it follows that $L_{z}$   has a quantised value $m\hslash$. Other quantum theory features such as the symmetry or anti-symmetry of the wave-function for identical particles are just assumed true in Bohmian mechanics. Relativistic versions of Bohmian mechanics are still being developed, as are quantum field theory versions. As far as is known, Bohmian mechanics does not predict any new physical results, but it has been applied to the well-known two-slit experiment for single particles [10] - where it is shown that the particle trajectories are affected by the quantum potential (that reflects the interference effects on the wave-function due to the two slits), and this then leads to the periodic particle position probability distribution on the observation screen. Whether Bohmian mechanics can describe bipartite systems for all pairs of non-commuting observables in each sub-system is an issue that is not examined in the present paper. However, interest in Bohmian mechanics and its extensions still continues [11].
1.5 Other Proposed Non Local HVT
Another approach to completing quantum theory is the Many Worlds Interpretation (MWI) [12]. Essentially the probabilistic feature of quantum theory - where in general each possible outcome of measuring an observable only occurs with a certain probability - is replaced with a deterministic picture in which all the different outcomes actually occur, but now in dfferent universes. The MWI has recently been discussed by Vaidman [13]. Leggett [14], introduced a so-called Crypto-Nonlocal (CN) hidden variable theory.In the context of the present paper, the LHVT expression (2) $P(\alpha|\Omega_{A},\lambda)P(\beta|\Omega_{B},\lambda)$ for the bipartite system case would be replaced in the CN theory by $P(\alpha|\Omega_{A},\Omega_{B},\lambda)P(\beta|\Omega_{A},\Omega_{B},\lambda)$, so that although the outcome for measuring either sub-system observable does not depend on the outcome for measuring the other sub-system observable, it does depend on the choice of the other observable. It is not clear whether this would produce different Bell inequalities from the NLHVT replacement $P(\alpha,\beta|\Omega_{A},\Omega_{B},\lambda)$. CN theory has been found to violate quantum theory and experiment [15].
@@ 1.6 Plan of Paper
First however, we need to outline the key equations describing both LHVT and a form of NLHVT in the simplest case of bipartite systems. Here the basic probability involves the outcomes for one observable in each sub-system. We then discuss the well-known CHSH example of a Bell inequality based on LHVT that is inconsistent with quantum theory. We also outline the idea behind LHVT. For reasons of completeness we next point out how Bell inequalities are also relevant to discussing quantum entanglement and EPR steering based on  a classification of Bell local states into three categories involving so-called local hidden quantum states [16]. The material outlining this is in Appendix A. We then introduce the so-called CGLMP Bell inequalities [7]. Here the basic probability involves the outcomes for two observables in each sub-system, resulting in more general LHVT andNLHVT forms. We then show a novel result - namely that the CHSH Bell inequality can also be proved for this more general NLHVT. The violation of the CHSH Bell inequality for this NLHVT approach proves that quantum theory cannot be underpinned by this type of NLHVT. The CGLMP inequalities themselves are applied in a bipartite system where each sub-system has two observables, each with the same number $d$ of outcomes) and set out the relevant equations for the inequalities - involving the quantity $S$, and introduce the observables and the test quantum state, both for the LHVT and NLHVT cases. The quantity $S$ involves the probabilities of all the outcomes for pairs of observables in the two sub-systems, where the outcomes are shifted by fixed amounts. The next section covers the basic matrix equation method for determining sufficiency requirements for possible Bell inequalities of the form $S\leq 4,S\leq 3,S\leq 2,S\leq 1$ in four cases for the NLHVT situation  Based on a consideration of all the sub-cases, sufficiency conditions for Bell tests in each of these four cases are set out, with the detailed derivations being presented in Appendix B. The number of sub-cases is $1,4,6,4$ for $S\leq 4,S\leq 3,S\leq 2,S\leq 1$ respectively. The conditions are in the form of inequalities involving the outcome shifts, and for the cases of $S\leq 2,S\leq 1$ , there are more than one such inequality for each of the sub-cases that are involved. At least one inequality must be satisfied for every sub-case. The last part of this section then sets out the quantum theory expressions for $S$. A comparison of the quantum theory numerical results for $S$ with the NLHVT results then follows. We again show that the CGLMP Bell inequality based on NLHVT is violated in quantum theory, reinforcing the previous conclusion based on CHSH. The paper ends with a Summary and Conclusions.
@@ 2 QTHY & HVT BASIC EQUATIONS
We consider the case of a bipartite system in which the two sub-systems may be spatialy separated. At  present we only consider one observable per sub-system- the  case of two will be considered later.
2.1 QTHY & HVT Probabilities - Bipartite Systems
The QTHY and LHVT joint probabilities $P(\alpha,\beta|\Omega_{A},\Omega_{B})$ for measurements of observables $\Omega_{A}$, $\Omega_{B}$ for sub-systems $A$, $B$ with outcomes $\alpha$, $\beta$ are:
@@ 2.1 QTHY & HVT Probabilities - Bipartite Systems
For QTHY $\widehat{\Pi}_{\alpha}^{A}$ is the projector onto the space of eigenvectors for operator $\widehat{\Omega}_{A}$ with eigenvalue $\alpha$. $\widehat{\rho}$ is density opr specifying the preparation process $c$ of the quantum state. For LHVT, hidden variables $\lambda$ determined with probability $P(\lambda|c)$ for preparation process $c$. The projectors satisfy $\sum_{\alpha}\widehat{\Pi}_{\alpha}^{A}=\widehat{1}^{A}$ and $\sum_{\beta}\widehat{\Pi}_{\beta}^{B}=\widehat{1}^{B}.$
In the probabilistic form of LHVT the hidden variables $\lambda$ determine each local sub-system measurement probabilities  $P(\alpha|\Omega_{A},\lambda)$, $P(\beta|\Omega_{B},\lambda)$ for classical observables $\Omega_{A}$, $\Omega_{B}$ which  are combined using classical probability theory to determine joint measurement probabilities. Since the probabilities for all outcomes must add up to unity we have $\sum_{\alpha}P(\alpha|\Omega_{A},\lambda)=1$ and $\sum_{\beta}P(\beta|\Omega_{B},\lambda)=1$. The origin of this expression for the LHVT probability expression involving the product of separate sub-system probabilities is set out below in Section 2.4. This form for the LHVT probability is widely used - apart from [5], see for example Refs. [16], [17], [9].
Both the QTHY and the LHVT probability satisfy the no-signaling condition. We have
$\sum_{\alpha}P(\alpha,\beta|\Omega_{A},\Omega_{B},c)_{Q}$ $=$ $Tr((\widehat{1}^{A}\otimes\widehat{\Pi}_{\beta}^{B})\widehat{\rho})=\sum_{\alpha^{\#}}P(\alpha^{\#},\beta|\Omega_{A}^{\#},\Omega_{B},c)_{Q}$
$\sum_{\beta}P(\alpha,\beta|\Omega_{A},\Omega_{B},c)_{Q}$ $=$ $Tr((\widehat{\Pi}_{\alpha}^{A}\otimes\widehat{1}^{B})\widehat{\rho})=\sum_{\beta^{\#}}P(\alpha,\beta^{\#}|\Omega_{A},\Omega_{B}^{\#},c)_{Q}$
$\sum_{\alpha}P(\alpha,\beta|\Omega_{A},\Omega_{B},c)_{LHVT}$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)P(\beta|\Omega_{B},\lambda)=\sum_{\alpha^{\#}}P(\alpha^{\#},\beta|\Omega_{A}^{\#},\Omega_{B},c)_{LHVT}$
$\sum_{\beta}P(\alpha,\beta|\Omega_{A},\Omega_{B},c)_{LHVT}$ $=$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)P(\alpha|\Omega_{A},\lambda)=\sum_{\beta^{\#}}P(\alpha,\beta^{\#}|\Omega_{A},\Omega_{B}^{\#},c)_{LHVT}$
where two different choices of sub-system observables $\Omega_{A}^{\#},\Omega_{B}^{\#}$ and outcomes $\alpha^{\#},\beta^{\#}$ are considered. Thus the outcome for sub-system’s observable is unaffected by the outcome or choice of the other sub-system’s observable.
If the LHVT form for the probability $P(\alpha,\beta|\Omega_{A},\Omega_{B},c)_{LHVT}$ does not apply, then what form would apply for NLHVT ? As explained in Sect. 2.4 the locality condition results in $P(\alpha,\beta|\Omega_{A},\Omega_{B},c)_{LHVT}$ involving the product of separate sub-system outcome probabilities as in Eq (2). Hence a general NLHVT probability will be one which  does not involve separate sub-system probabilities. Thus, in the probabilistic form of this general type of NLHVT, the hidden variables $\lambda$ should determine the combined sub-system measurement probabilities  $P(\alpha,\beta|\Omega_{A},\Omega_{B},\lambda)$, for both classical observables $\Omega_{A}$, $\Omega_{B}$  In this general non–local hidden variable theory there is no separate (or local) HVT probability for each sub-system.
@@
Again, the sum of the probabilities for all outcomes must be unity,so $\sum_{\alpha,\beta}P(\alpha,\beta|\Omega_{A},\Omega_{B},\lambda)=1$. This form for the NLHVT proability is the general form that would be involved when the local form in (2) is not applicable. Other more specific forms of NLHVT (such as Bohmian Mechanics) are briefly discussed in Sects. 1.4 and 1.5. There is no obvious reason why the no-signaling condition would apply in NLHVT.
@@ 2.2 Joint Measurement Mean Values
where $\left\langle\Omega_{A}(\lambda)\right\rangle=\sum_{\alpha}\alpha P(\alpha|\Omega_{A},\lambda)$ is LHVT mean value for measurement of $\Omega_{A}$ when hidden variables are $\lambda$. Similarly for $\left\langle\Omega_{B}(\lambda)\right\rangle$.
@@
2.3 CHSH Bell Inequality Test
@@
$|S|$ $=$ $|\left\langle\Omega_{A1}\otimes\Omega_{B1}\right\rangle_{LHVT}+\left\langle\Omega_{A1}\otimes\Omega_{B2}\right\rangle_{LHVT}+\left\langle\Omega_{A2}\otimes\Omega_{B1}\right\rangle_{LHVT}-\left\langle\Omega_{A2}\otimes\Omega_{B2}\right\rangle_{LHVT}\,|\,\leq 2$
Note the minus sign. This is violated for the singlet state of two spin $1/2$ systems (spin operators are in units /2Planck-constant-over-$\hbar/2$).
@@
for Pauli spin operators $\Omega_{A1}=\sigma_{z}^{A}$, $\Omega_{A2}=\sigma_{x}^{A}$, $\Omega_{B1}=-(\sigma_{x}^{B}+\sigma_{z}^{B})/\sqrt{2}$, $\Omega_{B2}=+(\sigma_{x}^{B}-\sigma_{z}^{B})/\sqrt{2}$, for which we find $S=2\sqrt{2}$ from quantum theory. This is based on the quantum theory mean value for the product of Pauli spin operator components along directions specified by unit vectors $\overrightarrow{a}$ and $\overrightarrow{b}$ given by $\left\langle\overrightarrow{\sigma}^{A}\cdot\overrightarrow{a}\otimes\overrightarrow{\sigma}^{B}\cdot\overrightarrow{b}\right\rangle_{Q}=-\overrightarrow{a}\cdot\overrightarrow{b}$ for the singlet state.
Note that the proof of the CHSH Bell inequality is based on the LHVT form (2) for the joint probability. It cannot be proved from the NLHVT form (4), though as we will see in Sect 3.2, the CHSH Bell inequality can be proved based on the two observable per sub-system NLHVT joint probability form (LABEL:Eq.NLHVTCGLMP).
@@
2.4 Origin of LHVT Expression
The origin of the LHVT joint probability expression (2) is explained in Ref [9]. Consider two events - $A$ being the measurement of observable $\Omega_{A}$ with outcome $\alpha$ and $B$ being the measurement of observable $\Omega_{B}$ with outcome $\beta$ , where in both cases when the hidden variables are $\lambda$. These events have a joint probability $P(B,A)$ for both occuring, whilst the probability of the separate events are $P(A)$, $P(B)$. Consider a situation (such as may occur when the events occur for separate widely separated sub-systems) where  the conditional probability $P(B|A)$ for event $B$ occuring given that event $A$ occurs has the same probability as if event $A$  does not occur. This is the situation of locality. Here the conditional probability $P(B|A)$ is independent of event $A$ and thus will not depend on the probability $P(A)$ of event $A$ occuring.
$P(B|A)$ $=$ $P(B)$ (10)
$P(B,A)$ $=$ $P(B|A)P(A)$ (11)
$=$ $P(B)\times P(A)=P(A)\times P(B)=P(A,B)$ (12)
after applying Bayes’ theorem. Thus the probability $P(B,A)=P(A,B)$ of both events occuring is equal to the product of the probabilities of each separate event occuring, and we obtain the LHVT expression (2)
$P(\alpha,\beta|\Omega_{A},\Omega_{B},\lambda)_{LHVT}$ $=$ $P(\alpha|\Omega_{A},\lambda)P(\beta|\Omega_{B},\lambda)$ (13)
after averaging over all hidden  variables $\lambda$- which occur with probability $P(\lambda|c)$. for preparation process $c$.
In the case where the hidden variable theory is non-local, the factorisation in Eq,.(13) does not apply (see Ref [16])  and $P(\alpha|\Omega_{A},\lambda)P(\beta|\Omega_{B},\lambda)$ would be replaced by $P(\alpha,\beta|\Omega_{A},\Omega_{B},\lambda)$, as in Eq (4).